Basic Scientific Computing

Marek Rychlik September 3, 2003

1 Booting

This section discusses the principles of bootoing a PC class machine. Let us try some equations:

A = πr^{2}  (1)

This is another equation.

∑_{n=1}^{∞} \frac{1}{n^{2}} = \frac{π^{2}}{6}

Yet another equation.

∫_{0}^{π} sin x dx = 0?